40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1349     Entry time: Mon Sep 23 11:57:13 2013
Author: Evan 
Type: DailyProgress 
Category: RF 
Subject: South PDH loop phase adjustment 

I removed 29 cm of SMA cable between the south RFPD RF output and the south TTFSS PD input in order to make the PDH error signal more symmetric. Relevant oscilloscope traces attached.

Attachment 1: southpdh.pdf  46 kB  | Hide | Hide all
southpdh.pdf
Attachment 2: data.zip  55 kB
ELOG V3.1.3-